A standing wave on a string that is fixed at both ends has frequency 80.0 Hz. The distance between adjacent antinodes of the standing wave is 16.0 cm. What is the speed of the waves on the string, in m/s?
Answer #1

antinodes are lambda/2 apart

so lambda = 32 cm

v = lambda f = .32*80= 25.6 m/s
Answer #2

speed of waves =distance* frequency =80*0.16=12.8 m/sec (ans)
Answer #3


lamda/2 = 16 cm...

lamda = 32 cm...= 0.32
speed v = lamda*f = 0.32*80 = 25.60 m/sec